write the slope-intercept equation for the line that passes through (-5,-8) and is perpendicular to 10x-6y=-11
Answer by mananth(16949)

passes through (-5,-8) and is perpendicular to 10x-6y=-11
10x-6y=-11
Find the slope of this line -6
y = -10x-11
Divide by -6
y=5/3 x+1.83
Compare this equation with y=mx+b
slope m =5/3

The slope of a line perpendicular to the above line will be the negative reciprocal
m1*m2=-1
The slope of the required line will be -0.60 -0.6

m= - 3/5 ,point ( -5 , -8 )
Find b by plugging the values of m & the point in
y=mx+b
-8 = 3.00 + b
b= -11
m= - 3/5
The required equation is y=- 3/5 x-11
m.ananth@hotmail.ca